Hi! I'm having trouble with part "c" on this word problem:

Hooke's Law states that the force needed to keep a spring stretched x units beyond its natural length is directly proportional to x.
(a) Write Hooke's Law as an equation. Use k for the constant of proportionality
(b) If a spring has a natural length of 8 cm and a force of F = 48 N is required to maintain the spring stretched to a length of 16 cm, find the spring constant
(c) What force is needed to keep the spring stretched to a length of 14 cm?
thanks!

Answer by scott8148(6628)

(a) F = k * x

(b) 48 = k * (16-8) ___ 48 = k * 8
dividing by 8 ___ 6 = k

(c) F = 6 * (14-8) ___ F = 6 * 6 ___ F = 36